Circular No. 8979

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 QP_{20} (LINEAR-HILL) R. E. Hill reports his discovery of a comet on Catalina Sky Survey un-

filtered CCD images taken with the 0.68-m Schmidt telescope, the object showing a well-condensed coma of diameter 6′′–8′′ and a narrow 18′′ tail in p.a. 250^{o} in four stacked 30-s exposures in 3′′-4′′ seeing (discovery observation tabulated below, though with magnitude added from Sept. 23.279 UT image); his stacked 60-s images from Sept. 24.3 show the comet to be well condensed with an elongated nuclear condensation with coma size 12′′ × 20′′ and a narrow 4′ tail in p.a. 260^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have reported this object to show cometary appearance, and at this point, K. Smalley (Minor Planet Center) identified Hill’s comet with an apparently asteroidal object discovered by LINEAR and designated 2008 QP_{20} on MPS 255263 (discovery observation also tabulated below). G. Muler (Lanzarate, Spain, 0.30-m Schmidt-Cassegrain reflector) writes that fifteen co-added 130-s exposures on Sept. 24.06 show a tail ∼ 29′′ long in p.a. 245^{o}.2. J. E. McGaha (Tucson, AZ, U.S.A., 0.62-m f/5.1 reflector) reports that a 180-s exposure from Sept. 24.2 shows a bright, round 7′′ coma with a narrow 44′′ tail in p.a. 243^{o}. H. Devore measured a 52′′.5 tail in p.a. 241^{o} from exposures taken by R. Holmes (Charleston, IL, U.S.A., 0.61-m f/4.0 astrograph) on Sept. 24.34–24.36. E. Guido, Castellammare di Stabia, Italy, notes that forty unfiltered co-added 60-s exposures taken remotely by G. Sostero, P. Camilleri, and himself (0.25-m f/3.4 reflector located near Mayhill, NM, U.S.A.) on Sept. 24.4 show a coma ∼ 10′′ in diameter and a narrow tail ∼ 30′′ long in p.a. 233^{o}. D. T. Durig (Sewanee, TN, U.S.A., 0.30-m f/2.5 Schmidt-Cassegrain reflector, Sept. 24.4, twenty 120-s co-added exposures) reports a tail ∼ 65′′ long in p.a. 245^{o}. A. C. Gilmore and P. M. Kilmartin (Mt. John 1.0-m f/7.7 reflector, Sept. 24.5, poor seeing) write that the object shows a stellar head and a tail ∼ 40′′ long in p.a. 240^{o}.

2008 UT α_{2000} δ_{2000} Mag. Observer
Aug. 25.39837 1^{h}35^{m}21.^{s}33 +18^{o}23′41′′.1 20.7 LINEAR
Sept. 23.26302 1 48 57.76 +24 38 26.1 17.7 Hill

The available astrometry, elliptical orbital elements (T = 2008 Nov. 2.277 TT, e = 0.50670, q = 1.72410 AU, P = 6.53 years, ω = 71^{o}.796, Ω = 325^{o}.167, i = 7^{o}.752, equinox 2000.0), and an ephemeris appear on MPEC 2008-S48.

2008 September 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green